# Google Tag Manager: Hidden Data Leaks and its Potential Violations under EU Data Protection Law

### 14 Dec 2023 v1

@@ 4.2.4. Tag Provider
S2: When the company behind the Tag Provider is not the same as the Data Collector, and the tag contains built-in consent, the Tag Provider includes the purposes for collecting data inside of the Tag’s code. This inclusion of predefined purposes in the Tag could qualify the Tag Provider as a Data Controller, since it seems to define the purposes itself. Even though the Tag Provider develops a data-collecting Tag by itself, which is released to the public and enables the collection of data, these facts are not enough to reveal a factual influence over the definition of purposes of personal data (AG-, 2023). We reason that by embedding the built-in consent in the tag, the Tag Provider might not relate to the processing of personal data in itself; it just consists of a prior step to the processing of personal data (AG-, 2023), it does not provide enough elements of a factual nature that the Tag Provider exercises an actual influence with regards to the “purposes and means” of processing data (AG-, 2023). Also, this actor does not process personal data. Hence, Tag Providers, under this scenario, are just service providers, with no GDPR obligations.
@@ 4.3.1. Server Tag Provider
This actor presents similar functions as to the Tag Provider on client side 4.2.4. On server-side GTM, the absence of the built-in function does not change its legal role as a Data Controller when the Server Tag Provider is also a Data Collector.

### 22 Dec 2023 v2

@@ 4.2.4. Tag Provider
S2: When the company behind the Tag Provider is not the same as the Data Collector, and the tag contains built-in consent, the Tag Provider includes the purposes for collecting data inside of the Tag’s code. This inclusion of predefined purposes in the Tag could qualify the Tag Provider as a Data Controller, since it seems to define the purposes itself. Even though the Tag Provider develops a data-collecting Tag by itself, which is released to the public and enables the collection of data, these facts are not enough to reveal a factual influence over the definition of purposes of personal data (AG-, 2023, par. 32). We reason that by embedding the built-in consent in the tag, the Tag Provider might not relate to the processing of personal data in itself; it just consists of a prior step to the processing of personal data (AG-, 2023), it does not provide enough elements of a factual nature that the Tag Provider exercises an actual influence with regards to the “purposes and means” of processing data (AG-, 2023). Also, this actor does not process personal data. Hence, Tag Providers, under this scenario, are just service providers, with no GDPR obligations.
@@ 4.3.1. Server Tag Provider
This actor presents similar functions as to the Tag Provider on client side (see §4.2.4). On server-side GTM, the absence of the built-in function does not change its legal role as a Data Controller when the Server Tag Provider is also a Data Collector.